# DataEnvGym: Data Generation Agents in Teacher Environments with Student Feedback

### v1

@@ 5 Conclusion
We propose DataEnvGym, a testbed of teacher environments for developing modular data generation agents (i.e., teachers) and environments. In DataEnvGym, a teacher agent generates training data for a student model and receives feedback on the student’s performance from the environment. We provide three environment instantiations with different levels of structure. Across three diverse domains (visual question answering, mathematics, and programming), we demonstrate that the example teachers we introduce can improve students in all DataEnvGym environments. We analyze DataEnvGym, including its training dynamics, performance across difficulty levels, and the impact of skill discovery module quality. We hope our work will foster future progress on data generation agents, engines, and feedback mechanisms by providing a testbed for evaluating and improving them.
@@ Acknowledgments
This work was supported by DARPA ECOLE Program No. HR00112390060, NSF-AI Engage Institute DRL-2112635, DARPA Machine Commonsense (MCS) Grant N66001-19-2-4031, ARO Award W911NF2110220, ONR Grant N00014-23-1-2356, Accelerate Foundation Models Research program, and a Bloomberg PhD Fellowship. The views contained in this article are those of the authors and not of the funding agency.
@@
In the appendix, we present additional related work (Appendix A), additional method details (Appendix B), and qualitative examples (Appendix C).
@@ B.3 Data Generation Details
B.4 Skill Forest Details
@@
B.5 Skill-Tree Policy
@@
B.6 Validation and Test Splits

### v2

@@ 5 Conclusion
We propose DataEnvGym, a testbed of teacher environments for developing modular data generation agents (i.e., teachers) and environments. In DataEnvGym, a teacher agent generates training data for a student model and receives feedback on the student’s performance from the environment. We provide three environment instantiations with different levels of structure. Across four diverse domains (visual question answering, mathematics, programming, and tool use), we demonstrate that the example teachers we introduce can improve students in all DataEnvGym environments. We analyze DataEnvGym, including its training dynamics, performance across difficulty levels, and the impact of skill discovery module quality. We hope our work will foster future progress on data generation agents, engines, and feedback mechanisms by providing a testbed for evaluating and improving them.
@@ Acknowledgments
This work was supported by DARPA ECOLE Program No. HR00112390060, NSF-AI Engage Institute DRL-2112635, DARPA Machine Commonsense (MCS) Grant N66001-19-2-4031, ARO Award W911NF2110220, ONR Grant N00014-23-1-2356, Microsoft Accelerate Foundation Models Research (AFMR) grant program, and a Bloomberg Data Science PhD Fellowship. The views contained in this article are those of the authors and not of the funding agency.
@@
In the appendix, we present additional related work (Appendix A), additional method details (Appendix B), and qualitative examples (Appendix C), experiments showing accuracy on generated data (Appendix D) and ablations on the relative importance of data vs training time (Appendix E).
@@ Appendix B Additional Method Details
In Fig. 6, we provide a more detailed walkthrough of the core environment loop. The example in Fig. 6 depicts one experience in the Open-Ended environment for the multimodal domain using the GQA dataset as the target dataset. The loop is highly modular, and can be changed to target a different data distribution by changing the dataset used for evaluation. The data generation policy and data generation engine can likewise be modified easily.
Figure 6: A complete data generation loop for the multimodal domain in the Open-Ended environment. Note that the teacher agent is modular. For domains such as mathematics or code generation, a data generation engine may not be required.
@@ B.3 Data Generation Details
B.4 Resource Costs
In Tab. 7, we tabulate the token costs and GPU time required for each environment and domain. The “Coding” domain and “Multimodal” domains are most expensive with respect to token expenditure and GPU time, respectively. Moreover, we show that we can reduce the overall cost by using a different teacher model. Specifically, we conduct experiments using GPT-4o-mini (which is less expensive than GPT-4o) as the LLM powering the data generation agent, which we tabulate in Tab. 6. Here, we continue to see performance improvements even with the reduced model. In general, we expect performance improvements to remain positive but decrease as we reduce the size of the teacher model.
Table 7: Token expenditure and GPU time per iteration for each environment and domain.Open-Ended environments are typically the cheapest. The most expensive domain is LiveCodeBench, although experiments in the Open-Ended for code cost less than \$3 per run. Training time for most environments, even with a single A6000 GPU, is typically less than 30 minutes.
Domain Environment Num Tokens Cost \$ (GPT-4o-mini) Cost \$ (GPT-4o) GPU Minutes / Iteration
Math Open-Ended 173,234 0.10 1.73 24
Math Skill-List 318,528 0.19 3.19 24
Math Skill-Tree 355,033 0.21 3.55 16
Coding Open-Ended 279,304 0.17 2.79 16
Coding Skill-List 497,787 0.30 4.98 16
Coding Skill-Tree 967,610 0.58 9.68 16
Multimodal Open-Ended 25,073 0.02 0.25 37
Multimodal Skill-List 82,419 0.05 0.82 134
Multimodal Skill-Tree 33,991 0.02 0.34 78
B.5 Skill Forest Details
@@
B.6 Skill-Tree Policy
@@
B.7 Validation and Test Splits
@@ Appendix C Qualitative Examples
Figure 16: Skill lists for each domain. These were determined by the skill discovery module and used in the Skill-Tree and Skill-List environments.
@@
Appendix D Accuracy on Generated Training Data
We analyze the performance of the student on training data generated by the teacher over successive iterations. We plot this data in Fig. 18. Concretely, for the Math and Multimodal environments, we plot the accuracy of a student at iteration $n$ on data generated in iteration $n+1$; thus, the data is generated but it has not been seen by the student yet. The student improves on data generated by the teacher through successive iterations as seen in Fig. 18. This shows that the data produced by the teacher is consistent over iterations and the student is gradually improving at the target data distribution. However, the data distribution produced by the teacher is complex enough that the improvement is neither monotonic or accomplished in a single iteration.
Figure 18: Over successive iterations, the student becomes more proficient at data generated by the teacher. Here, we plot the accuracy of a student from iteration $n$ on data generated by the teacher in iteration $n+1$, which has not been seen by the student.
Appendix E Ablation: Disentangling the effect of additional training from additional data
To substantiate that claim student performance is increased due to added data points rather than insufficient training, we take a subset of the data and increase the number of epochs such that the student receives a fraction of the added data, but an equivalent number of epochs as training on the full data. For example, if a student is normally trained for 10 epochs with 1000 generated training datapoints, we take the data from the first data generation iteration (which might, for example, contain 200 training datapoints) and train an alternative student for $\frac{1000}{200}\times 10=50$ epochs to isolate the effect of the generated training data vs the added training epochs. We show the results in Tab. 8. In each case, training with less data but for more epochs produces substantially smaller improvements than training with more data for fewer epochs, showing that data is responsible for increased performance rather than more training. In fact, extending training without additional data typically hurts performance — fresh data is essential. This highlights the importance of studying data generation as we do, as data generation is one of the few ways of obtaining fresh data.
Mutimodal (GQA) MATH Coding (LiveCodeBench)
Data Epochs Accuracy Data Epochs Accuracy Data Epochs Accuracy
Before Teaching - - 44.18 - - 15.78 - - 16.50
Less Data / Longer Training 20% 15 42.79 10% 30 13.98 20% 15 15.00
More Data / Standard Training 100% 3 47.90 100% 3 23.44 100% 3 18.91
Table 8: Training with less data but for more epochs produces significantly smaller improvements than training with more data for fewer epochs, showing that data is responsible for increased performance rather than more training.